OpenStudy (anonymous):

The measure of line XY is 48 and the coordinate of X is -5, What is the coordinate of y?

OpenStudy (anonymous):

we need the process :)

OpenStudy (anonymous):

oh wait you can just find the difference lol.

OpenStudy (anonymous):

d=y-x 48=y+5 43

OpenStudy (anonymous):

can we ask again? thank you so much

OpenStudy (anonymous):

the measure of line RS is 38 and the coordinate of R isa 14, what is the coordinate of S?